OpenStudy (anonymous):

how can i solve this equation: 1 1/3+ 3 3/4?

OpenStudy (anonymous):

turn mixed numbers into improper fractions and add them.

OpenStudy (anonymous):

4/3 + 15/4 4/3*4 = 16/12 15/4*3 = 45/12 16/12 + 45/12 = 61/12

OpenStudy (anonymous):

thank you guys

OpenStudy (nilankshi):

good ans